The Research Methods and Statistic module is almost over. Assume that the total marks were at an average of μ= 18 per assessment, and that the distribution of total marks are normally distributed with σ= 10.

What is the proportion that a student would have a mark more than 24 mark if randomly selected?


1
Expert's answer
2021-06-07T18:41:18-0400

Let "X=" the mark of student: "X\sim N(\mu, \sigma^2)"

Given "\mu=18, \sigma=10."

"P(X>24)=1-P(X\leq24)"

"=1-P(Z\leq\dfrac{24-18}{10})=1-P(Z\leq 0.6)"

"\approx0.247253"